OpenStudy (anonymous):

Write an equation in slope-intercept form of the line through point P(6, 7) with slope –2

OpenStudy (anonymous):

Write an equation in slope-intercept form of the line through point P(6, 7) with slope –2 y – 7 = –2(x – 6) y = –2x + 7 y = –2x + 19 y – 6 = –2(x – 7)

OpenStudy (anonymous):

y-7=-2(x-6) y-7=-2x+12 y=-2x+12+7 y=-2x+19

OpenStudy (anonymous):

witch one is the answer !!

OpenStudy (anonymous):

y=-2x+19 thats it

OpenStudy (anonymous):

y=-2x+19

OpenStudy (anonymous):

thankkkk you <333

OpenStudy (anonymous):

The first two are in point-slope form: y - y_1 = m(x - x_1) Therefore, it's not either of the first two. y = mx + b is slope-intercept form y = -2x + 19